Circular No. 5742

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 H. G. Corwin, California Institute of Technology, provides the following

UBV RI (R, I on the Cousins system; observers are requested to indicate if they are not using this system) photometry for stars GSC 4383.n in the vicinity of the supernova: n = 0224, V = 11.85, B − V = +1.30, U − B = +1.29, V − R = +0.68, V − I = +1.29; 0308, 11.78, +0.69, +0.16, +0.40, 0.80; 0434, 12.45, +0.63, +0.11, +0.39, +0.76; 0928, 11.90, +0.50, –0.08, +0.30, +0.60; 1023, 14.73, +0.46, +0.12:, +0.36, +0.60:; 1031, 11.62, +0.53, –0.04, +0.33, +0.69; 1123, 13.65, +0.57, –0.09:, +0.34, +0.69. Mean internal standard deviations are no more than ±0.03 mag, and none of the stars shows evidence of variability. The GSC V magnitudes of the seven stars are 11.17, 11.86, 12.00, 11.42, 14.36, 11.16 and 13.23, respectively.

A. C. Porter, Kitt Peak National Observatory, reports: “Harris V RIband CCD photometry by A. W. Neely, Silver City, NM, yields Mar. 31.21 UT, V = 10.33; Apr. 1.18, 10.60; 2.15, 10.97 (±0.03; comparison GSC 4383.0928 with GSC V). The R and I band photometry, not yet calibrated, indicates that the supernova is reddening, with the maximum occurring progressively later in redder bands: R faded by 0.07 and 0.25 and I by 0.04 and 0.16 mag during Mar. 31-Apr. 1 and during Apr. 1-2, respectively. Unfiltered magnitudes confirm that maximum occurred on about Mar. 31: Mar. 29.18 UT, 11.2; 31.21, 9.9; Apr. 1.18, 10.5; 2.15, 11.0. The Mar. 28 magnitude on IAUC 5740 was measured relative to GSC 4383.0340 (GSC V = 14.59). Preliminary Thuan-Gunn photometry from CCD images obtained by J. G. Cohen with the 1.5-m Palomar reflector on 1985 Jan. 21 shows the presumed progenitor at r = 21, g − r = +0.9. The measurements were made with a 5′′ aperture and 5′′-wide sky annulus to minimize background contamination. The uncertainty in the color may be as large as ±0.25 mag because of the galaxy background, but the redness is certain, and the value derived suggests a spectral type of ∼ M0.”

P. Prugniel, Observatoire de Haute Provence, reports that several 1min CCD images of the progenitor during Mar. 16.90-23.10 UT do not show evidence for variation. R(Cousins) = 20.0 ± 0.2. The object is at the detection limit in B and fainter than a star measured at B = 21.0 ± 0.3.

Photometry by D. Hanzl, Copernicus Observatory, Brno: Mar. 31.961 UT, V = 10.97 ± 0.04; B − V = +0.33 ± 0.03; Apr. 1.816, 11.19 ± 0.01, +0.32 ± 0.03 (comparison HD 86677, V = 7.876 ± 0.031, B − V = +0.510 ± 0.015). Single-channel UBV RI photometry by M. Mikolajewski and B. Wikierski, Copernicus University, Toruń: Apr. 1.86, V = 11.32, U − B = −0.56, B − V = +0.26, V − R = +0.37, V − I = +0.37.

1993 April 3 Brian G. Marsden